OpenStudy (anonymous):

Sin(cot+cos(tan))=cos+sin^2 an explanation would be helpful

OpenStudy (anonymous):

sorry, your supposed to prove it using iddentities

OpenStudy (amoodarya):

do you type it correct ?

hartnn (hartnn):

start by writing tan = sin/cos cot = cos/sin

OpenStudy (anonymous):

given that Sin(cot+cos(tan))=cos+sin^2 then Sin(cot+cos(tan)) can also writen as Sin(cot+cos(sin/cos)) then Sin(cot+sin) sin((cos/sin)+sin)= (sin(cos/sin)+sin*sin)= cos+sin^2

OpenStudy (anonymous):



OpenStudy (anonymous):

yes thankss

OpenStudy (anonymous):

Your are welcome......................